# NICE: To Optimize In-Context Examples or Not?

### v1

@@ A.2 Proposition Proofs and Discussion on the NICE Metric
In this section we prove that the NICE metric defined in Section 3 indeed follows the propositions in Section 3.1, and discuss some ways the NICE score can be used for prompt optimization.
@@
$\forall jS(j,\text{I},\mathcal{D})\approx\max\limits_{i}S(j,\text{I},\mathcal{D})$
@@
Let $b_{\text{max}}=\arg\max_{j}S(j,\text{I},\mathcal{D})$ be the best bin. If, for all $j\neq b_{\text{max}}$, $S(j,\text{I},\mathcal{D})\ll S(b_{\text{max}},\text{I},\mathcal{D})$, representing a strictly retrieval task where the budget needs to be spent on in-context example optimization. In this scenario: $\mathrm{\text{NICE}}\to 1/\lvert\mathcal{B}\rvert$ and $\lim\limits_{{\lvert\mathcal{B}\rvert\to\infty}}\mathrm{\text{NICE}}=0$.
@@
The bin-wise NICE metric can also be used for instruction/prompt optimization. One possible way is as follows: Start with an initial instruction that maximizes $\sum_{j}S(j,I)$. Then, using weights $w_{j}\propto S(j,I)$, optimize $\sum_{j}w_{j}S(j,I)$ to find the next instruction, and iterate. This is almost an alternating optimzation over instruction and examples (if not examples, the weights over different bins from which examples are chosen) that intuitively optimizes instructions for the examples (or bins) that matter more for the task at hand. We leave this exploration and other uses of NICE for future work.
A.3 NICE Results for Open-Source Models
NICE of various datasets with and without instructions for “Mixtral” (an 8x7B mixture-of-experts model) are shown in Fig. and for “Llama-2-70b-chat-hf” are shown in Fig. 12. We note similar observations as GPT-4 when using open-source language models.
Figure 11: Per-bin performance trend of various datasets with distance-based bin using “Mixtral”, a mixture-of-experts model. Arrow indicates change in NICE when shifting from no instruction to a detailed instruction. Note that the y-axis shows the ratio of performance of a given bin with the max performance over all bins with no instructions.
Figure 12: Per-bin performance trend of various datasets with distance-based bin using “Llama-2-70b-chat-hf”. Arrow indicates change in NICE when shifting from no instruction to a detailed instruction. Note that the y-axis shows the ratio of performance of a given bin with the max performance over all bins with no instructions.
A.4 Pre-trained Models: Is Instruction-tuning the key?
While the results demonstrated in Figs. 2, 12, and 11 validate our hypotheses for instruction-tuned models, we see in Fig. 13 that they are valid for pre-trained models as well. Pre-training on next-token prediction enables models to learn a probability distribution $p_{LM}(\mathbf{x})$ over variable length sequences, thus allowing them to learn task-specific priors empowering them with zero-shot learnability, as was observed by Choi et al. (2022). We see that even models not explicitly fine-tuned on instruction-following exhibit the ability to learn from contexts.
@@ A.4 Pre-trained Models: Is Instruction-tuning the key?
In Fig. 13, we show the results of the Llama-2-13B model on TREC (classification) and Break (generation) tasks. We see a partial pareto-dominance in case of semantic classification task like TREC (with high NICE) while it is not observed for the schema-based tasks like Break which has a low-NICE score.

### v3

@@ A.2 Proposition Proofs and Discussion on the NICE Metric
In this section we prove that the NICE metric defined in Section 4 indeed follows the propositions in Section 4.1, and discuss some ways the NICE metric can be used for prompt optimization.
@@
$\forall j\in\{1,...,\lvert B\rvert\},S(j,\text{I},\mathcal{D})\approx\max\limits_{i}S(j,\text{I},\mathcal{D})$
@@
Let $b_{\text{max}}=\arg\max_{j}S(j,\text{I},\mathcal{D})$ be the best bin. If, $\forall$ $j\neq b_{\text{max}}$, $S(j,\text{I},\mathcal{D})\ll S(b_{\text{max}},\text{I},\mathcal{D})$, representing a strictly retrieval task where the budget needs to be spent on in-context example optimization. In this scenario: $\mathrm{\text{NICE}}\to 1/\lvert\mathcal{B}\rvert$ and $\lim\limits_{{\lvert\mathcal{B}\rvert\to\infty}}\mathrm{\text{NICE}}=0$.
@@
A.2.1 Discussion on the NICE Metric
We can define a partial order on the set of instructions as $I\preceq I^{\prime}$ (in words, $I$ is pareto-dominated by another instruction $I^{\prime}$) if $S(j,I)\leq S(j,I^{\prime})$, $\forall j$. One can prove (under simple assumptions) that if there exists an $I^{*}$ that is not Pareto-dominated by any other $I^{\prime}$, then it must be a maximizer of $\sum_{j}w_{j}S(j,I)$ for some non-negative weights $w_{j}$’s (Harsanyi, 1955). The weights $w_{j}$’s are typically uniform or chosen by domain experts (Harsanyi, 1955; Gale, 1960). For simplicity, we used uniform weights to define our metric.
The bin-wise scores can also be used for instruction/prompt optimization. One possible way is as follows: Start with an initial instruction that maximizes $\sum_{j}S(j,I)$. Then, using weights $w_{j}\propto S(j,I)$, optimize $\sum_{j}w_{j}S(j,I)$ to find the next instruction, and iterate. This is almost an alternating optimzation over instruction and examples (if not examples, the weights over different bins from which examples are chosen) that intuitively optimizes instructions for the examples (or bins) that matter more for the task at hand. An optimized instruction should pareto-dominate all other possible instructions. We leave this exploration and other uses of NICE for future work.
A.3 Examples for the MTOP Dataset
We show some examples from the MTOP dataset in Fig. 11.
{promptbox}
Has Angelika Kratzer video messaged me?: \$[IN:GET_MESSAGE [SL:CONTACT Angelika Kratzer ] [SL:TYPE_CONTENT video ] [SL:RECIPIENT me ] ]\$
When will my next alarm start?: \$[IN:GET_ALARM [SL:ORDINAL next ] ]\$
Ich möchte gerne mit meinen Cousinen Ashlyn, Linda, Chloe, und Angel über Whatsapp telefonieren: \$[IN:CREATE_CALL [SL:CONTACT [IN:GET_CONTACT [SL:CONTACT_RELATED meinen ] [SL:TYPE_RELATION Cousinen ] [SL:CONTACT Ashlyn ] [SL:CONTACT Linda ] [SL:CONTACT Chloe ] [SL:CONTACT Angel ] ] ] [SL:NAME_APP Whatsapp ] ]\$
Figure 11: Some examples from the MTOP dataset.
A.4 NICE Results for Other Models
NICE-metric results for various datasets with and without instructions for GPT-3.5-Turbo are shown in Fig. 12 and for Mixtral-8x7B-v0.1 (an 8x7B mixture-of-experts model) and for Llama-2-70b-chat are shown in Figs. 13 and 14. We note similar observations as GPT-4-Turbo when using open-source language models.
Figure 12: Per-bin performance trend of various datasets with distance-based bin using GPT-3.5-Turbo. Arrow indicates change in NICE from no instruction to a detailed instruction.
Figure 13: Per-bin performance trend of various datasets with distance-based bin using Mixtral-8x7B-v0.1, a mixture-of-experts model. Arrow indicates change in NICE from no instruction to a detailed instruction.
Figure 14: Per-bin performance trend of various datasets with distance-based bin using Llama-2-70b-chat. Arrow indicates change in NICE from no instruction to a detailed instruction.
A.5 Pre-trained Models: Is Instruction-tuning the key?
While the results demonstrated in Figs. 2, 14, and 13 validate our hypotheses for instruction-tuned models, we see in Fig. 15 that they are valid for pre-trained models as well. Pre-training on next-token prediction enables models to learn a probability distribution $\text{p}_{\text{LM}}(\mathbf{x})$ over variable length sequences, thus allowing them to learn task-specific priors empowering them with zero-shot learnability, as was observed by Choi et al. (2022). We see that even models not explicitly fine-tuned on instruction-following exhibit the ability to learn from contexts.
@@ A.5 Pre-trained Models: Is Instruction-tuning the key?
In Fig. 15, we show the results of the Llama-2-13B model on TREC (classification) and Break (generation) tasks. We see a partial pareto-dominance in case of semantic classification task like TREC (with high NICE) while it is not observed for the schema-based tasks like Break which has a low NICE.
A.6 Additional Ablations and Results
A.6.1 Task Transformations
For further exploration into the kinds of tasks where ICE optimization matters, we designed the following ablation experiment: for the tasks with a standard output format (such as SST), we performed a transformation to the output space to induce a novel and unseen schema/structure to the output. Specifically, we defined the following mapping for the labels:
  • For SST2, we map:
    • negative $\rightarrow$ negative
    • positive $\rightarrow$ POSITIVE
  • For SST5, we map:
    • very negative $\rightarrow$ very negative
    • negative $\rightarrow$ negative
    • neutral $\rightarrow$ NeUtRaL
    • positive $\rightarrow$ POSITIVE
    • very positive $\rightarrow$ VERY POSITIVE
Since the new output labels are non-standard, we hypothesize that it would help to include similar in-context examples in the prompt, which are likely to share the same label, thus ensuring that the model is aware of the new label format. The observations from this ablation further strengthen our hypothesis.
The NICE scores with and without instructions in each case are given in Tab. 2, and the original scores can be found in Tab. 1.
Dataset NI DI
SST2 (original) 0.99 1.00
SST2 (induced schema) 0.86 0.96
SST5 (original) 0.91 0.94
SST5 (induced schema) 0.83 0.93
Table 2: NICE scores with and without instructions in original vs. induced schema on SST2 and SST5. Note the dip in performance without any instruction as we move to an induced schema.
Quantitatively, we observe the following:
  • A dip in the NICE score (corresponding to the No-Instructions case) as compared to standard datasets, thus showing the importance of learning the schema for model performance.
  • This dip (showing the dependence of ICE choice) recovers in presence of detailed instructions (since these tasks are easy enough to be learned from detailed task-specific instructions).
@@ A.6.1 Task Transformations
A.6.2 Correctness Metrics on GSM8k
We experiment with the GSM8k Cobbe et al. (2021) task to study the role of ICE optimization for complex mathematical reasoning tasks. We make use of 4-shot examples in each case. An example output for this dataset looks like this:
Natalia sold 48/2 = <<48/2=24>>24 clips in May. Natalia sold 48+24 = <<48+24=72>>72 clips altogether in April and May.
#### 72
We define two metrics to evaluate model outputs:
Final Answer Correctness (FAC) We consider the last numerical value in the string response to compare against the ground-truth and do not penalize the model for the format. The NICE scores with respect to the FAC metric are high ($0.91$ with no instructions to $0.98$ with detailed instructions), which means the task becomes almost invariant to choice of examples.
Reasoning Format Correctness (RFC): On the output string of the model, we use regular expressions to check if the model adheres to the reasoning syntax. In this case, we see that the NICE scores are $0.72$ (for no instructions) and $0.74$ (for detailed instructions), i.e., significantly lower than $0.91$/$0.98$, indicating that ICE optimization matters a lot for learning the correct format. This is consistent with our hypothesis about schema being the most important thing that is learned through examples.